### v1

@@ 1 Introduction
EDFs exist in different versions. Most well known are non-relativistic approaches that employ convenient representations of effective in-medium interactions between nucleons, e.g., the Skyrme or Gogny type models [1, 2, 3, 4]. Relativistic or covariant EDFs start from a field-theoretic approach where the interaction is described by meson-exchange taking medium dependent effects into account in different ways [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Either self-interactions of the mesons are introduced in so-called nonlinear (NL) models or a density dependence (DD) of the the meson-nucleon couplings is utilized. These extensions of the original Walecka model [16] are essential to obtain reasonable properties of nuclei and nuclear matter. Applying certain approximations to the full many-body wave function of the system of interest, e.g., of Hartree, Hartree-Fock or Hartree-Fock-Bogoliubov type, a set of coupled equations is found which has to be solved self-consistently.
@@
In the present work, new parameterisations of a relativistic EDF will be presented that include tensor couplings between nucleons and mesons in addition to the standard minimal couplings. Such an extension has been considered before in early applications of relativistic EDFs, see, e.g., Ref. [6], but it has been explored scarcely in depth. Including $\rho$-tensor, but no $\omega$-tensor, contributions in relativistic Hartree-Fock calculations showed that the single-nucleon spectrum could be improved [17]. The interest in relativistic EDFs with tensor couplings has been grown recently as it can affect the description of surface properties of nuclei and (indirectly) the equation of state of nuclear matter. New parametrisations were developed in Ref. [18] for relativistic models with meson-nucleon couplings that depend on vector or scalar densities of the medium in various combinations, but using a more limited number of nuclear observables in the parameter fit as compared to the present work. The effects of tensor couplings on different properties of nuclei and, using the resulting equation of state (EOS) of dense matter, on neutron stars were explored in Ref. [19], however without a thorough determination of model parameters. Here, we shall embark on the development of a new relativistic EDF with minimal and tensor couplings using a careful selection of nuclear observables that are used in the fit of the model parameters. The medium dependence of the interaction is taken into account with a density dependence of the minimal couplings using the functional forms introduced in Ref. [11].
@@ 2.1 Lagrangian density, field equations, and energy density
$\vec{\jmath}^{(t)}_{\omega}=\vec{\jmath}_{\pi}^{(t)}-\vec{\jmath}_{\nu}^{(t)}\>,\qquad\vec{\jmath}^{(t)}_{\rho}=\vec{\jmath}_{\pi}^{(t)}-\vec{\jmath}_{\nu}^{(t)}\>,$ (23)
@@ 2.2 Spherical nuclei
$\mathcal{Y}_{\kappa m}(\hat{r})=\sum_{m_{l}m_{s}}(l_{\kappa}\>m_{l}\>s\>m_{s}|j_{\kappa}\>m)Y_{l_{\kappa}m_{l}}(\hat{r})\chi_{sm_{s}}\>,$ (32)
in the upper and lower components of the Dirac spinor. Each state is identified by the index $\eta=\pi,\nu$, and two quantum numbers $n_{\eta}$, and $\kappa_{\eta}$, that are collectively denoted as $k_{\eta}$, and a total angular momentum projection $m_{\eta}$. The first, $n_{\eta}=0,1,2,\dots$, is the principal quantum number related to the number of nodes of the function $F_{\eta k_{\eta}}$; the second, $\kappa_{\eta}=\pm 1,\pm 2,\dots$, defines the total angular momentum $j_{\kappa}=|\kappa|-1/2$ and the orbital angular momentum $l_{\kappa}=j_{\kappa}-\kappa/(2|\kappa|)$ of a state. Then the Dirac equation (10) is equivalent to the set of two coupled differential equations;
$\left(E_{\eta k_{\eta}}-V_{\eta}-M_{\eta}+S_{\eta}\right)F_{\eta k_{\eta}}+\left(\frac{d}{dr}+\frac{\kappa_{k}}{r}-T_{\eta}\right)G_{\eta k_{\eta}}=0\>,$ (33)
@@
$\left(\frac{d}{dr}-\frac{\kappa_{k}}{r}+T_{\eta}\right)F_{\eta k_{\eta}}-\left(E_{\eta k_{\eta}}-V_{\eta}+M_{\eta}-S_{\eta}\right)G_{\eta k_{\eta}}=0\>,$ (34)
@@ 3.2.1 Values of observables from theory
of the single-nucleon energies of the levels with identical principal quantum number $n_{\eta}$, and $k_{\eta}^{\pm}=(n_{\eta},\kappa_{\eta}^{\pm})$, where $\kappa_{\eta}^{\pm}=(j_{\eta}^{\pm}-l_{\eta})(2j_{\eta}^{\pm}+1)$ for $j_{i\eta}^{\pm}=l_{\eta}\pm 1/2$. The constraint isoscalar monopole giant resonance energy $E_{\mathrm{mono}}$ is obtained from a constraint calculation of the nucleus as described in Ref. [1]. A contribution
@@ 3.2.2 Selection of nuclei and experimental data
The description of nuclei with the relativistic EDF presented here can only be applied to nuclei with spherical symmetry. Also pairing effects are not taken into account. Thus a comparison of theory and experiment can be reasonably performed only for magic or semi-magic nuclei. This limits the amount of data that can be included in the fit of the model parameters. However, this restriction is not too severe since the essential features of strongly interacting systems are sufficiently well constrained by the selection of the fitted data in this approach. In order to fix the isospin dependence of the effective interaction well, nuclei with large $|N-Z|$ should be included, preferably on both sides of the valley of stability. The nuclei actually considered can be read off from Tables 6, and 7. There are 18 (semi)magic nuclei and 2 with one valence neutron above a doubly magic core. Not all data from the list in Subsection 3.2 are always available for all nuclei.
@@ 3.4 Types of energy density functionals
The main aim of the present work is the development of a new parameterisation of a relativistic EDF with tensor couplings, denoted by ’DDT’ in the following, in addition to the usual minimal couplings of the mesons to the nucleons. In order to compare the main differences of this EDF to a model without tensor interactions, the results of the DD2 model [34] will be shown, that uses the same form of the relativistic EDF.
@@ 4 Results and discussion
The correlations coefficients of the fit parameters in the DDT model are given in Figure 1 where boxes in red color indicate the largest absolute values whereas grey boxes correspond to the smallest moduli of the coefficients. As expected, a first group with $M_{\sigma}$, $M_{\mathrm{nuc}}^{\ast}$, and $\Gamma_{T\omega}$ shows strong correlations since they all affect the description of the nuclear surface. In contrast, $\Gamma_{T\rho}$ is hardly correlated to any other parameter of the fit. The other two parameters related to the isovector part of the interaction, $J$, and $L$, however, show a strong correlation. Apart from this dependence, $J$ seems to be practically uncorrelated to all other fit parametes. Another pair of parameters, $r_{10}$, and $r_{21}$, determines the density dependence of the isoscalar couplings. They are also strongly correlated. A somewhat lesser strength of correlation is found for the incompressibility coefficient $K$, and $r_{10}$ with the first group of parameters.
@@ 4.2 Nuclear matter
$E_{0}(n_{b})=-B_{\mathrm{sat}}+\frac{1}{2}Kx^{2}+\frac{1}{6}Qx^{3}+\dots\>,$ (91)
and skewness parameter
$E_{\mathrm{sym}}(n_{b})=J+Lx+\frac{1}{2}K_{\mathrm{sym}}x^{2}+\dots\>,$ (94)
and the symmetry incompressibility
$K_{\tau,v}=K_{\rm sym}-6L-\frac{QL}{K}\>,$ (97)
that describes the change of the incompressibility with the variation in $\alpha$ [15, 46]. These quantities, together with the effective Dirac mass at saturation $(M_{\mathrm{nuc}}^{\ast})_{\mathrm{sat}}=M_{\mathrm{nuc}}-C_{\sigma}n_{\mathrm{sat}}^{(s)}$, are convenient to characterize the equation of state of nuclear matter around saturation. They are given in Table 12 for the models DDT, DDTC, and DD2. Most remarkable are the increased effective mass and the reduced incompressibility coefficient of the models with tensor couplings (DDT and DDTC) in comparison to that without these couplings (DD2). Most existing relativistic models show values of $K$ that are rather high when compared with non-relativistic EDFs, cf. Refs. [15, 3]. The tensor models are an exception and they are also characterized by rather large saturation densities. The higher-order coefficients $Q$ and $K_{\mathrm{sym}}$ show considerable differences and are not very well determined by the parameter fits as they have large uncertainties. The smaller values of $K$ and $Q$ of the DDTC model indicate a softer equation of state as compared to the DDT model as seen already in Figure 5. The DDT and DDTC models also have symmetry energy slope parameters that are smaller than the value of $L$ of the DD2 model. This difference correlates with the smaller neutron skin thicknesses for exotic nuclei as discussed in Subsection 4.1.
@@
$\left(M_{\rm nuc}^{\ast}\right)_{\mathrm{sat}}$ $M_{\mathrm{nuc}}$ $0.65729$ $(0.01315)$ $0.66305$ $(0.03803)$ $0.56252$
@@ 5 Summary and Conclusions
The EDFs with the new parameterisations can be studied in the future for larger sets of nuclei, including deformed ones, taking other effects like pairing into account. The additional tensor couplings, however, require some extra work to be considered in such deformed calculations. The new models can also be used to prepare EOS tables of stellar matter at finite temperature for larger ranges in isospin asymmetry and density, in order to be used in simulations of core-collapse supernovae and neutron-star mergers [57]. For this, the calculations have to be extended by including additional degrees of freedom, e.g., light clusters at subsaturation densities [34, 58], or hyperons at high densities, see, e.g., Ref. [59].
@@ Appendix A Transformation between parameters of nuclear matter and of coupling functions
    The reference mass $\varrho_{\mathrm{ref}}$ is chosen as the baryon density at saturation $n_{\mathrm{sat}}$ and defines the Fermi momentum

### v2

@@ 1 Introduction
EDFs exist in different versions. Most well known are non-relativistic approaches that employ convenient representations of effective in-medium interactions between nucleons, e.g., the Skyrme or Gogny type models [1, 2, 3, 4]. Relativistic or covariant EDFs start from a field-theoretic approach where the interaction is described by meson-exchange taking medium dependent effects into account in different ways [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Either self-interactions of the mesons are introduced in so-called nonlinear (NL) models or a density dependence (DD) of the the meson-nucleon couplings is utilized. Another class of relativistic EDFs uses only nucleons as degrees of freedom with point couplings and strengths that are constants or depend on the density, see, e.g., Refs. [16, 17, 18, 19]. No explicit meson contributions appear in these formulations. These extensions of the original Walecka model [20] are essential to obtain reasonable properties of nuclei and nuclear matter. Applying certain approximations to the full many-body wave function of the system of interest, e.g., of Hartree, Hartree-Fock or Hartree-Fock-Bogoliubov type, a set of coupled equations is found which has to be solved self-consistently.
@@
In the present work, new parameterisations of a relativistic EDF will be presented that include tensor couplings between nucleons and mesons in addition to the standard minimal couplings. Such an extension has been considered before in early applications of relativistic EDFs, see, e.g., Ref. [6], but it has been explored scarcely in depth. Effects of considering $\omega$-tensor couplings in EDFs on the spin-orbit splittings and indirectly on the nucleon effective mass have been studied in Ref.  [21]. A relativistic EDF with $\omega$ and $\rho$-tensor couplings in the framework of the NL coupling models has been developed in Ref. [22] taking into account also contributions from Dirac-sea nucleons. The parameters of this model were adjusted to the energies of a eight nuclei from 16O to 208Pb. Including $\rho$-tensor, but no $\omega$-tensor, contributions in relativistic Hartree-Fock calculations showed that the single-nucleon spectrum could be improved [23]. Tensor couplings were also included in the framework of point couplings models in the Hartree-Fock approximation, see, e.g., Ref. [24]. The interest in relativistic EDFs with tensor couplings has been grown recently as it can affect the description of surface properties of nuclei and (indirectly) the equation of state of nuclear matter. New parametrisations were developed in Ref. [25] for relativistic models with meson-nucleon couplings that depend on vector or scalar densities of the medium in various combinations, but using a more limited number of nuclear observables in the parameter fit as compared to the present work. A new parameterisation of a covariant EDF with density dependent minimal couplings and density independent tensor couplings in the Hartree-Bogoliubov approximation was developed recently [26]. The model parameters were determined using constraints from nuclear matter and nuclei with preset uncertainties of the fitted observables. The effects of tensor couplings on different properties of nuclei and, using the resulting equation of state (EOS) of dense matter, on neutron stars were explored in Ref. [27], however without a thorough determination of model parameters. Here, we shall embark on the development of a new relativistic EDF with minimal and tensor couplings using a careful selection of nuclear observables that are used in the fit of the model parameters. Particular attention is paid to the choice of the uncertainties of the considered observables. The medium dependence of the interaction is taken into account with a density dependence of the minimal couplings using the functional forms introduced in Ref. [11]. In addition to the description of nuclei, the implications of the new parameterisations on the equation of state of nuclear matter and the properties of neutron stars will be explored.
@@ 2.1 Lagrangian density, field equations, and energy density
$\vec{\jmath}^{(t)}_{\omega}=\vec{\jmath}_{\pi}^{(t)}+\vec{\jmath}_{\nu}^{(t)}\>,\qquad\vec{\jmath}^{(t)}_{\rho}=\vec{\jmath}_{\pi}^{(t)}-\vec{\jmath}_{\nu}^{(t)}\>,$ (23)
@@ 2.2 Spherical nuclei
$\mathcal{Y}_{\kappa_{\eta}m}(\hat{r})=\sum_{m_{l}m_{s}}(l_{\eta}\>m_{l}\>s\>m_{s}|j_{\eta}\>m_{\eta})Y_{l_{\eta}m_{l}}(\hat{r})\chi_{sm_{s}}\>,$ (32)
in the upper and lower components of the Dirac spinor. Each state is identified by the index $\eta=\pi,\nu$, and two quantum numbers $n_{\eta}$, and $\kappa_{\eta}$, that are collectively denoted as $k_{\eta}$, and a total angular momentum projection $m_{\eta}$. The first, $n_{\eta}=0,1,2,\dots$, is the principal quantum number related to the number of nodes of the function $F_{\eta k_{\eta}}$; the second, $\kappa_{\eta}=\pm 1,\pm 2,\dots$, defines the total angular momentum $j_{\eta}=|\kappa_{\eta}|-1/2$ and the orbital angular momentum $l_{\eta}=j_{\eta}-\kappa_{\eta}/(2|\kappa_{\eta}|)$ of a state. Then the Dirac equation (10) is equivalent to the set of two coupled differential equations;
$\left(E_{\eta k_{\eta}}-V_{\eta}-M_{\eta}+S_{\eta}\right)F_{\eta k_{\eta}}+\left(\frac{d}{dr}+\frac{\kappa_{\eta}}{r}-T_{\eta}\right)G_{\eta k_{\eta}}=0\>,$ (33)
@@
$\left(\frac{d}{dr}-\frac{\kappa_{\eta}}{r}+T_{\eta}\right)F_{\eta k_{\eta}}-\left(E_{\eta k_{\eta}}-V_{\eta}+M_{\eta}-S_{\eta}\right)G_{\eta k_{\eta}}=0\>,$ (34)
@@ 2.3 Nuclear matter
$E_{0}(n_{b})=\frac{E}{A}(n_{b},0)=-B_{\mathrm{sat}}+\frac{1}{2}Kx^{2}+\frac{1}{6}Qx^{3}+\dots\>,$ (60)
and skewness parameter
$E_{\mathrm{sym}}(n_{b})=J+Lx+\frac{1}{2}K_{\mathrm{sym}}x^{2}+\dots\>,$ (63)
and the symmetry incompressibility
$K_{\tau,v}=K_{\rm sym}-6L-\frac{QL}{K}\>,$ (66)
that describes the change of the incompressibility with the variation in $\alpha$ [15, 30]. These quantities, together with the effective Dirac mass at saturation $M_{\mathrm{sat}}^{\ast}=M_{\mathrm{nuc}}^{\ast}(n_{\mathrm{sat}})=M_{\mathrm{nuc}}-C_{\sigma}n_{\mathrm{sat}}^{(s)}$, are convenient to characterize the equation of state of nuclear matter around saturation.
@@ 3.2.1 Values of observables from theory
of the single-nucleon energies of the levels with identical principal quantum number $n_{\eta}$, and $k_{\eta}^{\pm}=(n_{\eta},\kappa_{\eta}^{\pm})$, where $\kappa_{\eta}^{\pm}=(j_{\eta}^{\pm}-l_{\eta})(2j_{\eta}^{\pm}+1)$ for $j_{\eta}^{\pm}=l_{\eta}\pm 1/2$. The constraint isoscalar monopole giant resonance energy $E_{\mathrm{mono}}$ is obtained from a constraint calculation of the nucleus as described in Ref. [1]. A contribution
@@ 3.2.2 Selection of nuclei and experimental data
The description of nuclei with the relativistic EDF presented here can only be applied to nuclei with spherical symmetry. Also pairing effects are not taken into account. Thus a comparison of theory and experiment can be reasonably performed only for magic or semi-magic nuclei. This limits the amount of data that can be included in the fit of the model parameters. However, this restriction is not too severe since the essential features of strongly interacting systems are sufficiently well constrained by the selection of the fitted data in this approach. In order to fix the isospin dependence of the effective interaction well, nuclei with large $|N-Z|$ should be included, preferably on both sides of the valley of stability. The nuclei actually considered can be read off from Tables 6, 7, and 9. There are 18 (semi)magic nuclei and 2 with one valence neutron above a doubly magic core. Not all data from the list in Subsection 3.2 are always available for all nuclei.
@@ 3.4 Types of energy density functionals
The main aim of the present work is the development of a new parameterisation of a relativistic EDF with tensor couplings, denoted by ’DDT’ in the following, in addition to the usual minimal couplings of the mesons to the nucleons. In order to compare the main differences of this EDF to a model without tensor interactions, the results of the DD2 model [43] will be shown, that uses the same form of the relativistic EDF. Although a different fitting strategy with smaller numbers of nuclei and observables was used to determine the parameters of the DD2 model, a comparison of the results with those of models with tensor couplings will be instructive even though the differences will originate not completely from these additional contributions in the EDF.
@@ 4 Results and discussion
The correlation coefficients of the fit parameters in the DDT model are given in Figure 1 where boxes in red color indicate the largest absolute values whereas grey boxes correspond to the smallest moduli of the coefficients. As expected, a first group with $M_{\sigma}$, $M_{\mathrm{sat}}^{\ast}$, and $\Gamma_{T\omega}$ shows strong correlations since they all affect the description of the nuclear surface. In contrast, $\Gamma_{T\rho}$ is hardly correlated to any other parameter of the fit. The other two parameters related to the isovector part of the interaction, $J$, and $L$, however, show a strong correlation. Apart from this dependence, $J$ seems to be practically uncorrelated to all other fit parametes. Another pair of parameters, $r_{10}$, and $r_{21}$, determines the density dependence of the isoscalar couplings. They are also strongly correlated. A somewhat lesser strength of correlation is found for the incompressibility coefficient $K$, and $r_{10}$ with the first group of parameters.
@@ 4.1 Nuclei
The contributions with tensor couplings in the EDF modify the strength of the spin-orbit interaction as compared to models without these couplings. Thus a modification in the single-particle level structure can be expected. This could also affect the appearance of shell gaps. As an example, the semi-magic proton shell gaps in 140Ce and 218U at $Z=58$ and $Z=92$, respectively, are examined here. It has been observed that many relativistic EDFs predict large spurious shell gaps for these nuclei, see, e.g., Refs. [47, 48, 24]. In the DD2 model the $\pi 0g_{7/2}-\pi 1d_{5/2}$ splitting in 140Ce amounts to $3.054$ MeV and the $\pi 0h_{9/2}-\pi 1f_{7/2}$ splitting in 218U is $3.248$ MeV. In fact, reduced values of $2.288$ MeV ($2.543$ MeV) and $2.665$ MeV ($2.999$ MeV) are found for the $Z=58$ and $Z=92$ proton shell gaps in the DDT (DDTC) model.
@@ 4.2 Nuclear matter
Characteristic nuclear matter parameters are given in Table 12 for the models DDT, DDTC, and DD2. Most remarkable are the increased effective mass and the reduced incompressibility coefficient of the models with tensor couplings (DDT and DDTC) in comparison to that without these couplings (DD2). Most existing relativistic models show values of $K$ that are rather high when compared with non-relativistic EDFs, cf. Refs. [15, 3]. The tensor models are an exception and they are also characterized by rather large saturation densities. The higher-order coefficients $Q$ and $K_{\mathrm{sym}}$ show considerable differences and are not very well determined by the parameter fits as they have large uncertainties. The smaller values of $K$ and $Q$ of the DDTC model indicate a softer equation of state as compared to the DDT model as seen already in Figure 5. The DDT and DDTC models also have symmetry energy slope parameters that are smaller than the value of $L$ of the DD2 model. This difference correlates with the smaller neutron skin thicknesses for exotic nuclei as discussed in Subsection 4.1.
@@
$M_{\mathrm{sat}}^{\ast}$ $M_{\mathrm{nuc}}$ $0.65729$ $(0.01315)$ $0.66305$ $(0.03803)$ $0.56252$
@@ 5 Summary and Conclusions
EDFs with the new parameterisations can be applied in further calculations to explore their quality in the description of finite nuclei. The study of deformed nuclei is of particular interest as it allows to investigate additional observables, e.g., higher multipole moments of the density distributions, or the variation of energies of single-particle states in isotopic and isotonic chains of nuclei. The tensor couplings could affect the position of the levels in the spectrum and give rise to the appearance of new shell gaps or their suppression. Effects of pairing have to be included in the model too when open-shell nuclei are considered. However, additional work with proper numerical methods is required to include deformation and pairing in the calculations. This is beyond the scope of the present work. The new models can also be used to prepare EOS tables of stellar matter at finite temperature for larger ranges in isospin asymmetry and density, in order to be used in simulations of core-collapse supernovae and neutron-star mergers [67]. For this, the calculations have to be extended by including additional degrees of freedom, e.g., light clusters at subsaturation densities [43, 68], or hyperons at high densities, see, e.g., Ref. [69].
@@ Acknowledgments
Statements and Declarations
Open Access funding enabled and organized by Projekt DEAL. The authors have no relevant financial or non-financial interests to disclose. Data sets generated during the current study are available from the corresponding author on reasonable request.
@@ Appendix A Transformation between parameters of nuclear matter and of coupling functions
    The reference density $\varrho_{\mathrm{ref}}$ is chosen as the baryon density at saturation $n_{\mathrm{sat}}$ and defines the Fermi momentum